# A comprehensive theory for relativistic polaritonic chemistry: a way toward a four component ab initio treatment of molecular systems

### v1

@@ IV Conclusions
In this paper we proposed a reformulation of relativistic QED allowing for an easier development of ab initio methodologies to simulate heavy atoms molecular complexes in strong coupling conditions. Using this theoretical ground, we reported the development and implementation of the first relativistic polaritonic wave function based ab initio method, namely Pol-DHF. The theory has been presented starting from the standard Lagrangian and has been derived into a usable implementation of the Pol-DHF code. Considering the possible competition with the polaritonic effects, the radiative QED corrections have been explicitly included into the treatment. After providing our roadmap to the implementation, we presented applications of Pol-DHF to three metal hydrides: CuH, AgH and, AuH. These systems were an excellent test case in order to assess the magnitude of the polaritonic effects in comparison to the relativistic effects. To do so, we evaluated the influence of the polaritonic effects on the gap for the three systems. The polaritonic effects resulted having the largest relative influence on AuH for which relativistic effects are more prominent. Afterwards, we provided a detailed analysis of the competition between the Gaunt, Breit and polaritonic effects on the ground state and orbital energies. We have shown that even though the Full-Breit term represents a significantly larger contribution to the total ground state energy, its effect is extremely orbital dependent. In particular, it is much larger on the core orbitals compare to the valence, where instead the polaritonic effects dominate. Therefore, we could conclude that neglecting the Full-Breit term (practice usually applied in relativistic quantum chemistry) in strong coupling conditions is only possible if properties involving valence orbitals need to be investigated. Lastly, we presented excited state calculations for AuH at the TD-Pol-DHF level. As already observed by Koneckny et al. [30], we have shown that using a fully relativistic polaritonic theory the appearance of Rabi-Splittings at the crossing between singlet and triplet potential energy surfaces can be observed. These results clearly open new possibilities for the control of the inter-system crossing in photochemistry. We strongly believe that the methodologies and the applications presented in this paper will represent a groundbreaking step toward the simulation of relativistic molecular systems strongly coupled to photons, field that is recently finding many interesting applications in photochemistry but also in spintronics and quantum computation.
@@ Acknowledgements.
Guillaume Thiam, Riccardo Rossi and Enrico Ronca acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon Europe Research and Innovation Program (Grant ERC-StG-2021-101040197-QED-SPIN). Henrik Koch acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation Program (Grant Agreement 101020016).Appendix A Length gauge transformation
@@
It is important to emphasize that the dipole approximation strongly simplifies the expression of the terms involving the momentum. The other noticeable fact is that the terms involving cancels out (the translated momentum bring out a term cancelling out with the unmodified term).

### v2

@@ IV Conclusions
In this paper, we proposed a reformulation of relativistic QED allowing for an easier development of ab initio methodologies to simulate heavy atoms molecular complexes in strong coupling conditions.
Using this theoretical ground, we reported the development and implementation of the first relativistic polaritonic wave function based ab initio method, namely Pol-DHF. The theory has been presented starting from the standard Lagrangian and has been derived into a usable implementation of the Pol-DHF code. Considering the possible competition with the polaritonic effects, the importance of radiative QED corrections have been addressed. After providing our road-map to the implementation, we presented applications of Pol-DHF to three metal hydrides: CuH, AgH and, AuH. These systems were an excellent test case in order to assess the magnitude of the polaritonic effects in comparison to the relativistic effects. To do so, we evaluated the influence of the polaritonic effects on the gap for the three systems. The polaritonic effects resulted having the largest relative influence on AuH for which relativistic effects are more prominent.
Afterward, we provided a detailed analysis of the competition between the Gaunt, Breit and polaritonic effects on the ground state and orbital energies.
We have shown that even though the full-Breit term represents a significantly larger contribution to the total ground state energy, its effect is extremely orbital dependent.
In particular, its impact is much stronger on the core orbitals compared to the valence, where instead the polaritonic effects dominate. Therefore, we could conclude that neglecting the full-Breit term (a common practice in relativistic quantum chemistry) in strong coupling conditions is only possible if properties involving exclusively valence orbitals need to be investigated. A similar study could be conducted adding effective potentials to include radiative QED corrections and comparing the magnitude of their effects to the polaritonic ones, this will surely be the topic of a future work.
Lastly, we presented excited state calculations for AuH at the TD-Pol-DHF level. As already observed by Konecny et al. [35], we have shown that using a fully relativistic polaritonic theory, the appearance of Rabi-Splittings at the crossing between singlet and triplet potential energy surfaces can be observed. These results clearly open new possibilities for the control of the inter-system crossing in photochemistry.
We strongly believe that the methodologies and the applications presented in this paper can represent a groundbreaking step toward the simulation of relativistic molecular systems strongly coupled to photons, field that is recently finding many interesting applications in photochemistry but also in spintronics and quantum computation.
@@ Acknowledgements.
The authors acknowledge Michael Ruggenthaler, Lukas Konecny, Rosario Roberto Riso and Giovanni Bistoni for insightful discussions. G. T., R. R. and E. R. acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon Europe Research and Innovation Program (Grant ERC-StG-2021-101040197-QED-SPIN). H. K. acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation Program (Grant Agreement 101020016).Appendix A Length gauge transformation
@@
It is important to emphasize that the dipole approximation strongly simplifies the expression of the terms involving the momentum. The other noticeable fact is that the terms involving cancels out (the translated momentum bring out a term canceling out with the unmodified term).